OpenStudy (anonymous):

use Logarthmic Differentiation to solve: y=(sin(3x))^ln(x)

OpenStudy (anonymous):

twig can u go back to my ? about jrotc

myininaya (myininaya):

take natural log of both sides \[\ln(y)=\ln[(\sin(3x))^{\ln(x)}]\] using properties of log learned in algebra we can write \[\ln(y)=\ln(x) \ln(\sin(3x))\] now take derivative of both sides \[\frac{y'}{y}=[\ln(x)]'\ln(\sin(3x)]+[\ln(x)][\ln(\sin(3x)]'\]

myininaya (myininaya):

let me know if you need help with the rest

OpenStudy (anonymous):

HEY JAMES

OpenStudy (anonymous):

Twig are you a girl!!

OpenStudy (anonymous):

myininaya I FINALLY got it thanks to you!!! Thanks a bunch! and bieberfan, yes I am

myininaya (myininaya):

np